40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Apr 1 12:01:52 2019, Jon, Update, CDS,  
    Reply  Mon Apr 1 22:33:00 2019, gautam, Update, CDS, ITMY freed 
       Reply  Tue Apr 2 15:02:53 2019, Jon, Update, CDS, ITMY freed 
Message ID: 14506     Entry time: Mon Apr 1 22:33:00 2019     In reply to: 14505     Reply to this: 14508
Author: gautam 
Type: Update 
Category: CDS 
Subject: ITMY freed 

While Anjali is working on the 1um MZ setup, the pesky ITMY was liberated from the OSEMs. The "algorithm" :

  • Apply a large (-30000 cts) offset to the side coil using the fast system.
  • Approach the zero of the YAW DoF from -2.00V, PIT from +10V (you'll have to jiggle the offsets until the optic is free swinging, and then step the bias down by 0.1). At this point I had the damping off.
  • Once the PIT bias slider reaches -4V, I engaged all damping loops, and brought the optic to its nominal bias position under damping. 

While doing this work, I noticed several errors corresponding to EPICS channel conflicts. Turns out the c1susaux2 EPICS server was left running, and the MEDM screens (and possibly several scripts) were confused. There has to be some other way of testing the new crate, on an isolated network or something - please do not leave the modbus service running as it potentially interferes with normal IFO operation. For good measure, I stopped the process and shut down the machine since I saw nothing in the elog about any running tests.

Quote:

ITMY became stuck during this process

ELOG V3.1.3-